Given a random variable X which is normally distributed with mean 15 and Standard deviation 10 find?

(a). P( X<20)


(b). P(X>12)


(c). P(12


1
Expert's answer
2021-09-15T02:57:22-0400

(a) "P(X<20)=P(Z<\frac{20-15}{10})=P(Z<0.5)=0.6915."


(b)

"P(X>12)=P(Z>\frac{12-15}{10})=P(Z>-0.3)=1-P(Z<-0.3)=\\=1-0.3821=0.6179."


(c) "P(12<X<20)=P(X<20)-P(X<12)=0.6915-0.3821=0.3094."